If the sum of 2 numbers are 40, and their difference is 6, what are these 2 numbers?
Answer by checkley77(12844)

X+Y=40
X-Y=6 ADD.
------------
2X=46
X=46/2
X=23 ANS.
23+Y=40
Y=40-23
Y=17 ANS.